OpenStudy (anonymous):

7 and fifteen-sixteenths minus 4 and seven-sixteenths is equal to what number?

OpenStudy (zzr0ck3r):

7+(15/16)-4-(7/16)

OpenStudy (johnweldon1993):

\[7\frac{ 15 }{ 16 } - 4\frac{ 7 }{ 16 }\] since these have the same denominator...we can simply subtract the whole numbers...and subtract the fractions 7 - 4 =...? (15/16) - (7/16)= ...?

OpenStudy (anonymous):

7 and one-fifth minus 2 and three-tenths is equal to what number?

OpenStudy (johnweldon1993):

For that one we can change the mixed numbers into fractions \[7\frac{ 1 }{ 5 } = \frac{ 36 }{ 5 }\] and \[2\frac{ 3 }{ 10 } = \frac{ 23 }{ 10 }\] \[\frac{36}{5} - \frac{23}{10} \] what is the least common denominator here?

OpenStudy (anonymous):

7 and fifteen-sixteenths minus 4 and seven-sixteenths is equal to what number?